STRONG VIBRONIC COUPLING IN THE VISIBLE SPECTRUM OF ZIRCONIUM METHYLIDYNE

\underline{ANTHONY J. MERER} and JAMES R.D. PEERS^{a}, Department of Chemistry, University of British Columbia, 2036 Main Mall, Vancouver, BC, Canada V6T 1Z1; SCOTT J. RIXON, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC, Canada V6T 1Z1.

Zirconium methylidyne, ZrCH, can be prepared by the reaction of methane (8% in helium) with laser-ablated zirconium atoms. At least 50 vibrational bands have been observed in the 565 − 665 nm region for each of ZrCH and ZrCD using laser-induced fluorescence. Dispersed emission spectra have allowed the assignment of numerous excited vibrational levels in the X̃^{2}Σ^{+} ground state (17 for ZrCH, 22 for ZrCD and 9 for Zr^{13}CH). The ground state bond lengths, as determined from high-resolution spectra of ZrCH and ZrCD, are r_{0}(Zr−C) = 1.83067(13) Å and r_{0}(C−H)= 1.0866(9) Å. P′ values derived from rotational analyses of 20 subbands in the electronic spectrum (9 of ZrCH and 11 of ZrCD) reveal a very complicated level structure. Preliminary vibrational assignment suggests the presence of two interacting excited states, 2Π and ^{2}Δ, about 250 cm^{−1} apart. Many vibronically-induced bands consequently appear with unusual properties such as anomalous branch intensity patterns or upper levels with very large electrostatic (Renner-Teller) splittings between their angular momentum components.

aPresent address: Creo Inc., 3755 Willingdon Avenue, Burnaby, BC, Canada V5G 3H3